Gail is willing to pay $89 for a new pair of shoes and Karen is willing to pY $60. What is the gain in total consumer surplus if the price of the shoes falls from $70 to $50?

A. $10

B. $19

C. $29

D. $30
Answer #1

Initial consumer surplus = (89-70) = 19

Now karen will also buy as price is now less than his willingness to pay.

New consumer surplus = (89-50) + (60-50) = 49

Increase in consumer surplus = 49-19= 30

D. $30